OpenStudy (anonymous):

What is the value of the expression 4^3 • 4^−5

OpenStudy (kirbykirby):

\(\Large a^b \cdot a^c=a^{b+c}\), so pretend here that \(a\) is 4, \(b\) is 3 and \(c\) is -5

OpenStudy (anonymous):

Add the two exponents because you are multiplying and then multiply like usual \[4^{3}\times 4^{(-5)}= 16^{3+(-5)}=16^{(-2)}\]